# A sufficient condition for generalized spectral characterization of graphs with loops

### v1

@@ 1. Introduction
A significant open problems in the field is Haemers’ conjecture, which posits that almost all graphs are characterized by their adjacency spectrum [16, 6]. Unfortunately, it is generally difficult to determine whether a given graph is characterized by its adjacency spectrum. While there are many methods that one can try to use to produce another graph with the same spectrum [14, 1, 4, 19], there is no easy way to verify that a graph to be characterized by its adjacency spectrum aside from exhaustive enumeration, unless the graph belongs to some special class whose structure can be exploited [3, 15, 7].
@@
While Theorem 1.2 also allows simple graphs, it is vacuous in this case. Indeed, it can be shown that $2^{\lfloor n/2\rfloor}$ always divides the determinant of the walk matrix for simple graphs [17, Lemma 14]. For graphs with loops, however, it can occur that $2^{2}$ is not a divisor of the determinant; see Figure 1 for an example.
@@
Adding loops hence has the nice feature that it eliminates the need to treat the prime $p=2$ separately. This was the main motivation for the current work, as this makes it less complicated to understand how often the condition is applicable. In particular, random matrix models that are expected to behave similarly to the adjacency matrix of a random graph with loops are analyzed in a forthcoming joint work of Lvov and the present author [10] leading to an explicit conjectural values for the satisfaction frequency of Theorem 1.2. Comparable random matrix analysis is expected to be more complicated for Theorem 1.1 due to the difficulties that surround the prime $2$ in that case.
@@
We establish a result in a more general framework. Consider an vector $\zeta\in\mathbb{Z}^{n}$ and an integral matrix $\mathbf{X}\in\mathbb{Z}^{n\times n}$ that is symmetric $\mathbf{X}=\mathbf{X}^{\top}$. Two pairs $(\mathbf{X},\zeta)$ and $(\mathbf{Y},\eta)$ are isomorphic up to sign if there exists a signed permutation $\mathbf{S}$ such that $\mathbf{S}\mathbf{X}\mathbf{S}^{\top}=\mathbf{Y}$ and $\mathbf{S}\zeta=\eta$. A signed permutation is a matrix of the form $\mathbf{S}=\mathbf{P}\mathbf{D}$ with $\mathbf{P}\in\{0,1\}^{n\times n}$ a permutation and $\mathbf{D}$ a diagonal matrix with diagonal entries $\pm 1$.
@@ Acknowledgement
Funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy EXC 2044 –390685587, Mathematics Münster: Dynamics–Geometry–Structure.

### v2

@@ 1. Introduction
A significant open problems in the field is Haemers’ conjecture, which posits that almost all graphs are characterized by their adjacency spectrum [16, 6]. Unfortunately, it is generally difficult to determine whether a given graph is characterized by its adjacency spectrum. While there are many methods that one can try to use to produce another graph with the same spectrum [14, 1, 4, 19], there is no easy way to verify that a graph is characterized by its adjacency spectrum aside from exhaustive enumeration, unless the graph belongs to some special class whose structure can be exploited [3, 15, 7].
@@
While Theorem 1.2 also allows simple graphs, it is vacuous in this case. Indeed, it can be shown that $2^{\lfloor n/2\rfloor}$ always divides the determinant of the walk matrix for simple graphs [17, Lemma 14]. For graphs with loops, it can occur that $2^{2}$ does not divide the determinant; see Figure 1 for example. Adding loops hence has the nice feature that it eliminates the need to treat the prime $p=2$ separately.
@@
Our proofs also enable generalizations of Theorem˜1.2 where the adjacency matrix may be replaced by an arbitrary symmetric integer matrices and the all-ones vector by an arbitrary integer vector; see Theorem˜2.12. Aside from its independent interest, this general setting is helpful to simplify probabilistic analysis, which was our main motivation for the present paper. In particular, random matrix models that are expected to behave similarly to graphs with loops are analyzed in a forthcoming work of Lvov and the present author [10], leading to conjectures for the satisfaction frequency of Theorem˜2.12 when the integer vector is also taken to be random. We return to this point in Section˜3.
@@
We establish a result in a more general framework. Consider a vector $\zeta\in\mathbb{Z}^{n}$ and an integral matrix $\mathbf{X}\in\mathbb{Z}^{n\times n}$ that is symmetric $\mathbf{X}=\mathbf{X}^{\top}$. Two pairs $(\mathbf{X},\zeta)$ and $(\mathbf{Y},\eta)$ are isomorphic up to sign if there exists a signed permutation $\mathbf{S}$ such that $\mathbf{S}\mathbf{X}\mathbf{S}^{\top}=\mathbf{Y}$ and $\mathbf{S}\zeta=\eta$. A signed permutation is a matrix of the form $\mathbf{S}=\mathbf{P}\mathbf{D}$ with $\mathbf{P}\in\{0,1\}^{n\times n}$ a permutation and $\mathbf{D}$ a diagonal matrix with diagonal entries $\pm 1$.
@@ 2.4. Proof of the sufficient condition
3. Conjecture for the satisfaction frequency
As mentioned in the introduction, Theorem˜2.12 is convenient for probabilistic investigations, which was our main motivation for the present paper. In particular, analytically tractable random matrix models are analyzed in our forthcoming work [10]. We believe that those models are in the same universality class as graphs with loops, leading to the following Conjecture 3.1:
Conjecture 3.1 (Lvov and Van Werde [10]).
Suppose that $\mathbf{X}$ is chosen uniformly at random from all symmetric $n\times n$ matrices with $\{0,1\}$-valued entries and consider an independent random vector $\zeta$ that is uniform on $\{0,1\}^{n}$. Then,
$\lim_{n\to\infty}\mathbb{P}\bigl(\det(\mathbf{W}_{\mathbf{X},\zeta})\textnormal{ is square-free}\bigr)$ $=\prod_{p\textnormal{ prime}}\Bigl(1-\frac{1}{p^{2}}-\frac{1}{p^{3}}+\frac{1}{p^{4}}\Bigr)\prod_{k=1}^{\infty}\Bigl(1-\frac{1}{p^{2k}}\Bigr)$
$=0.2943\ldots$ (3.1)
Similar conjectures are expected to be more difficult for graphs without loops, or if $\zeta$ is taken to be the all-ones vector. In particular, additional complications are necessary in that setting related to special behavior for the prime $p=2$.
@@ Acknowledgement
I thank Nikita Lvov for discussions related to the topic of this paper. Funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy EXC 2044 –390685587, Mathematics Münster: Dynamics–Geometry–Structure.